Grade 9Mechanics

A mine worker measures his weight inside a mine and find that it has decreased by 0.05% of that on surface of earth. Then find the depth of mine

Maaruf
7 Years agoGrade 9

1 Answer

Khimraj
7 Years ago
g’ = 99.95*g/100
g(1 – d/R) = 99.95*g/100
1- d/R = 1- 0.05/100
d/R = 0.05/100
d = 0.05*6400/100 = 3.2 KM
Hope it clears.............................